# Restricted Receptive Fields for Face Verification

### v1

@@ IV-B Verification Rates for RRFNet
While competitive performance is achieved with $56\times 56$ patches, a notable performance drop for cross-pose datasets is observed for $28\times 28$ patches in Tab. II when using region-based similarities. We hypothesize that this decline primarily arises from restricting comparisons to corresponding patches at the same spatial positions. To overcome this limitation, we introduce a second approach, RRFNet, which compares each patch in one image with all patches in the other image to compute a comprehensive similarity score. For instance, RRFNet-28 divides each $112\times 112$ face image into 33 patches of size $28\times 28$, resulting in $33\times 33=1089$ patch pairs used to measure image-level similarity (illustrated in Fig. 2).
TABLE IV: Four patch configurations for $56\times 56$ receptive fields. The configuration in the last row is chosen for the experiments as it achieves the highest accuracy.
@@
To determine the receptive fields for RRFNet-56, we evaluate four different configurations for the $56\times 56$ setting, as shown in Tab. IV. The highest verification rates are obtained using 5 patches located at $(0,28)$, $(28,0)$, $(56,28)$, $(28,56)$, and $(28,28)$ in our preliminarily experiments. To balance accuracy and computational cost, we do not consider configurations with more patches, and this 5-patch setup is used to report results. Note, this choice excludes the four $28\times 28$ corner regions of the $112\times 112$ face images as shown in Fig 3. For RRFNet-28, which uses $28\times 28$ patches, the same positions are applied, yielding a total of 33 patches. Compared to the baseline approach [15] with $112\times 112$ input, this results in approximately a $3\times$ increase in training time for RRFNet-56 and a $7\times$ increase for RRFNet-28. Note, although our approach is slower than the ResNet, it does not require any additional computation to generate an explanation of the similarity decision as opposed to post-hoc approaches.
Verification performance of the proposed approach can be seen in Tab. III. Publicly-available models are used for a comparison with three recent approaches [14, 66, 67]. As we use the same architecture and the loss function, we trained models for the ArcFace [15] in the same settings to report results for a fair comparison. Our RRFNet-28 and RRFNet-56 models trained for 30 epochs. During training [5,5][-5,5] vertical and horizontal shifts are applied as data augmentation. Also, effect of masked patch augmentation is analyzed with two masked patch ratio $20\%$ and $40\%$. Two datasets, Casia-WebFace [68] and WebFace4M [64], and two CNN architectures, ResNet50 and ResNet100 [57], are used for training recognition models using the publicly-available implementation [65].
@@
As shown in Tab. III, RRFNet-56 consistently achieves the highest verification rates across 7 datasets under different training settings. The only exception is CPLFW, where KP-RPE [67] reports $95.40$ accuracy. Note that while KP-RPE requires an additional model for keypoint supervision, our architecture does not rely on any positional information and processes all patches uniformly. Surprisingly, we observe that RRFNet-28 achieves competitive results, even though verification is performed using patches as small as $28\times 28$. These results demonstrate the effectiveness of the patch-level similarity approach while providing more insight into the composition of face similarity.
V Conclusion
In this work, we propose a face similarity metric using patch-level similarity scores. Compared to the traditional approach, using a single feature vector representing the entire face image, we extract multiple representations from restricted receptive fields. These representations are used to calculate patch similarity scores between two images. Then, combination of patch similarities are used to express the global similarity to make a binary decision. In contrast to using a holistic face representation for measuring the similarity between images, our approach brings inherent explanation to the decision process as the similarity between faces is decomposed into local similarities.
@@ V Conclusion
Unlike post-hoc explanation methods, which require additional processing after a decision is made, our approach adopts a simple design choice to improve model transparency without requiring extra computation to generate an explanation. In addition to its interpretable properties, it achieves competitive verification rates using patches as small as $28\times 28$. Moreover, we show the proposed RRFNet-56 achieves higher verification rates than the state-of-the-art approaches.
@@
In future work, we will investigate the automatic selection of receptive field sizes and positions to further enhance performance. Although our uniformly distributed patches across the face image yield strong results, we believe that adapting patch selection to the specific image pairs under comparison could further improve verification rates. Moreover, while we strictly follow the configurations of [15] to ensure a fair performance comparison, we believe our approach could be further optimized using alternative CNN or ViT architectures.

### v2

@@ IV-B Verification Rates for RRFNet
While competitive performance is achieved with $56\times 56$ patches, a notable performance drop for cross-pose datasets is observed for $28\times 28$ patches in Table II when using region-based similarities. We hypothesize that this primarily arises from restricting comparisons to corresponding patches at the same spatial positions. To overcome this limitation, we introduce a second approach, RRFNet, which compares each patch in one image with all patches in the other image to compute a comprehensive similarity score. For instance, RRFNet-28 divides each $112\times 112$ face image into 33 patches of size $28\times 28$, resulting in $33\times 33=1089$ patch pairs used to measure image-level similarity.
To determine the receptive fields for RRFNet-56, we evaluate four different configurations for the $56\times 56$ setting, as shown in Table V. The highest verification rates are obtained using 5 patches located at $(0,28)$, $(28,0)$, $(56,28)$, $(28,56)$, and $(28,28)$ in our preliminarily experiments. To balance accuracy and computational cost, we do not consider configurations with more patches, and this 5-patch setup is used to report results. Note, this choice excludes the four $28\times 28$ corner regions of the $112\times 112$ face images as shown in Fig. 2. For RRFNet-28, which uses $28\times 28$ patches, the same positions are applied, yielding a total of 33 patches.
Verification performance of the proposed approach is given in Table III. Publicly-available models of three recent approaches [14, 69, 79] are used for a comparison. As we use the same architecture and the loss function, we trained models for the ArcFace [15] in the same settings to report results for a fair comparison. Our RRFNet-28 and RRFNet-56 models trained for 30 epochs. During training [5,5][-5,5] vertical and horizontal shifts are applied as data augmentation. Also, effect of masked patch augmentation is analyzed with two masked patch ratio $20\%$ and $40\%$. Two datasets, Casia-WebFace [80] and WebFace4M [77], and two CNN architectures, ResNet50 and ResNet100 [60], are used for training recognition models using the publicly-available implementation [78].
TABLE V: Four patch configurations for $56\times 56$ receptive fields. The configuration in the last row is chosen for the experiments as it achieves the highest accuracy.
@@
TABLE VI: Comparison with explainable face verification approaches.
Method # Param. Training Time (h) Inference Time (s) LFW CFPFP CPLFW
Ins. $(\uparrow)$ Del. $(\downarrow)$ Ins. $(\uparrow)$ Del. $(\downarrow)$ Ins. $(\uparrow)$ Del. $(\downarrow)$
ArcFace [15]
(w/ random saliency)
$65M$ $17$ $0.11$
$0.658$
$\pm 0.062$
$0.658$
$\pm 0.060$
$0.601$
$\pm 0.084$
$0.600$
$\pm 0.084$
$0.616$
$\pm 0.109$
$0.617$
$\pm 0.109$
ArcFace [15]
(w/ xSSAB [22])
$65M$ $17$ $0.52$
$0.660$
$\pm 0.071$
$0.481$
$\pm 0.093$
$0.588$
$\pm 0.112$
$0.476$
$\pm 0.121$
$0.635$
$\pm 0.129$
$0.466$
$\pm 0.130$
ArcFace [15]
(w/ xFace [23])
$65M$ $17$ $61$
$0.787$
$\pm 0.042$
$0.282$
$\pm 0.058$
$0.805$
$\pm 0.055$
$0.262$
$\pm 0.092$
$0.761$
$\pm 0.101$
$0.263$
$\pm 0.089$
xCos [56]
(Self-interpretable)
$103M$ 12 $0.09$
$0.647$
$\pm 0.071$
$0.250$
$\pm 0.060$
$0.532$
$\pm 0.118$
$0.354$
$\pm 0.136$
$0.544$
$\pm 0.122$
$0.348$
$\pm 0.137$
RRFNet-28
(Self-interpretable)
$56M$ $81$ $0.29$
$0.822$
$\pm 0.028$
$0.199$
$\pm 0.032$
$0.852$
$\pm 0.031$
$0.166$
$\pm 0.043$
$0.829$
$\pm 0.069$
$0.186$
$\pm 0.072$
RRFNet-56
(Self-interpretable)
$65M$ $42$ $0.21$
$0.819$
$\pm 0.031$
$0.223$
$\pm 0.035$
$0.829$
$\pm 0.036$
$0.208$
$\pm 0.044$
$0.797$
$\pm 0.095$
$0.234$
$\pm 0.087$
Reference  Probe
0% mask 5% mask 10% mask 15% mask
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
score=0.62 score=0.49 score=0.33 score=0.21
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
score=0.45 score=0.24 score=0.21 score=0.18
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
score=0.44 score=0.48 score=0.49 score=0.48
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
score=0.43 score=0.45 score=0.50 score=0.50
Figure 4: Examples of impostor (top two) and genuine (bottom two) pairs on Hadrian dataset. Similarity scores are given below each probe image. For impostor pairs, salient regions are masked, while for genuine pairs least salient regions are masked.
Figure 5: FPR and FNR analysis on Hadrian dataset.
@@
As shown in Table III, RRFNet-56 consistently achieves the highest verification rates across 7 datasets under different training settings. The only exception is CPLFW, where KP-RPE [79] obtains $95.40$ accuracy. While KP-RPE requires an additional model for keypoint supervision, our architecture does not rely on any positional information and processes all patches uniformly. Surprisingly, we observe that RRFNet-28 achieves competitive results, even though verification is performed using patches as small as $28\times 28$. Note, the verification rates achieved with $28\times 28$ patches show a significant improvement over the region-based similarity approach (see Table II comb. $28\times 28$). To further investigate this gain, we measure the percentage of patches whose most similar match in genuine pairs are located at the same spatial position. Results are reported in Table IV for RRFNet-28. We observe that this ratio can drop to as low as $20\%$ in cross-pose scenarios, highlighting the importance of cross-region patch matching compared to the region-based similarity approach.
V Analysis on Explainability
RRFNet-28 xFace
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Examples of genuine image pairs. For each probe image, the saliency map and corresponding insertion and deletion scores are shown for both RRFNet-28 and xFace [23].
In this section, we compare our approach with prior works on explainability methods for face verification. Two post-hoc approaches, xSSAB [22] and xFace [23], and a self-interpretable method, xCos [56], are used for the comparison. For post-hoc methods, ArcFace [15] is used as the backbone network.
Evaluation Metrics. Faithfulness of the generated saliency maps is measured using the Insertion and Deletion metrics [81]. The Insertion Area Under the Curve (AUC) score quantifies the increase in similarity scores as salient facial regions are progressively revealed, whereas the Deletion AUC score measures the decrease in similarity as salient regions are removed. In addition, the False Positive Rate (FPR) and False Negative Rate (FNR) are used to analyze the impact of masking the most salient region in impostor pairs and the least salient region in genuine pairs.
Comparison with previous approaches is presented in Table VI. Genuine pairs from the LFW, CFPFP, and CPLFW datasets are used to report Insertion and Deletion scores for evaluating the quality of saliency maps. Training time is measured using four L40S GPUs, and inference time is reported for a pair of images on a CPU. We observe that, our approach outperforms previous works in both Insertion and Deletion metrics. Although xFace [23] achieves scores closest to ours, its inference time is substantially higher, as it requires 2,260 forward passes to generate a single saliency map. Examples for comparison between RRFNet-28 and xFace are shown in Fig. 6. In some cases, the Insertion and Deletion scores of xFace fall to the level of random chance, highlighting the vulnerability of perturbation-based post-hoc methods.
Next, we analyze the effect of masking salient and least salient regions on impostor and genuine pairs, respectively. FPR and FNR curves for different masking rates are shown in Fig. 5, and examples are presented in Fig. 4. Masking is performed by applying Gaussian blurring. We observe that the most salient regions typically correspond to the nose and eye areas, and masking these regions leads to a sharp decrease in similarity scores for impostors. Note, while a similar impact is also observed for genuine pairs, we investigate the impact of masking least salient regions to observe whether such masking can increase similarity scores. Our analysis shows that, while some pairs exhibit increased similarity after masking, typically on hair region, the overall FNR remains unchanged.
VI Conclusion
@@
In this work, we propose RRFNet to enable a joint optimization of local feature learning and score fusion, defining global face similarity as a sum of local similarities. We demonstrate that this approach enables competitive verification rates using patches as small as small as $28\times 28$. Moreover, we show that RRFNet-56 can achieve higher verification rates than the state-of-the-art approaches. Unlike post-hoc explanation methods that require additional processing after a decision is made, our approach improves model transparency through its inherent design, eliminating the need for extra computation to generate explanations.
@@
In future work, we plan to investigate the automatic selection of receptive field sizes and positions to further enhance performance. Although our uniformly distributed patches across the face image yield strong results, we believe that adapting patch selection to the specific image pairs under comparison could further improve verification rates. Moreover, while we strictly follow the configurations of [15] to ensure a fair performance comparison, we believe our approach could be further optimized using alternative CNN or ViT architectures.
Acknowledgment
Kagan Ozturk is partly supported by the Ministry of National Education of Türkiye.